OpenStudy (anonymous):

tom is 8 years older than jane.Five years ago Tom was three times as old as Jane. How old is Jane now?

OpenStudy (anonymous):

Because you need an equation where there is an 8 year difference and where you can multiply the smaller number by 3 you get the larger. 12=4*3 because 12-4=8 and 4*3 =12 so you know Jane was 4, 5 years ago. So now, 5 years later, Jane is 4+5=9